# thanks: These three authors contributed equallythanks: These three authors contributed equallythanks: These three authors contributed equally

### v1

@@ Abstract
I Introduction
Topological order [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18] is a profound concept in condensed matter physics, describing phases of matter that cannot be characterized by traditional symmetry-breaking mechanisms [19]. Topological degeneracy is one of the most fundamental aspects of systems exhibiting topological order. Topological degeneracy refers to the ground state degeneracy of a system with topological order, which grows exponentially with the genus of the manifold on which the system is defined [20, 21, 22].
@@ I Introduction
In contrast to the toric code, systems with more complex topological order typically exhibit non-Abelian exchange phases [4, 8], indicating that their ground states are strongly-correlated and cannot be described within a quasi-particle framework. This strong correlation makes the construction of such topologically ordered systems on a lattice particularly challenging. However, in this work, we demonstrate that flat bands can emerge in systems exhibiting topological degeneracy without necessarily requiring non-Abelian exchange phases.
@@
The paper is organized as follows. In Section II, we introduce the Lieb lattice as an example of a singular flat band. We also demonstrate how the number of NLEs increases with the genus of the lattice and establish the relationship between the number of holes and zero-energy states under different boundary conditions. In Section III, we calculate the entanglement entropy of a quasi-one-dimensional Lieb ring. In Section IV, we investigate the square lattice as an example of a perfectly nested van Hove system and propose a method for constructing flat bands at the van Hove energy. Section V concludes with a brief summary of our results and a discussion of the implications of our findings.
II flat band from a singular flat band system
II.1 Lieb lattice on a 1-genus surface
The Hamiltonian of the Lieb lattice[68, 69, 76, 77], considering only nearest-neighbor hopping, is given by:
@@ II.1 Lieb lattice on a 1-genus surface
II.2 Lieb lattice on a 2-genus surface
As introduced previously, the linearly independent NLEs of a Lieb lattice with periodic boundary conditions can be geometrically interpreted as non-contractible loops of the torus. As a result, when the Lieb lattice is placed on a 2-genus surface, there are four linearly independent NLEs, as seen in Fig.  3(b). To validate this, we diagonalize the Hamiltonian in Eq. 1 with $\mu\neq 0$ to obtain all the zero-energy states, which collectively form the Hilbert space $V_{E=0}$. By determining the Hilbert space of all CLEs, denoted as $V_{CLE}$, the number of linearly independent NLEs, $N_{NLE}$, can be expressed as:
@@ II.2 Lieb lattice on a 2-genus surface
II.3 Lieb lattice with defects
The topological degeneracy of the Lieb lattice can also be manifested in an open-boundary Lieb lattice with multiple hole defects. Similar to the toric code [29], the Lieb lattice can be punctured in two physically distinct ways, leading to two different types of boundary conditions. These boundary conditions are crucial in shaping the spatial configurations of CLE and NLE. As illustrated in Fig. 4(a) and (c), removing five lattice sites to create a hole causes four CLEs to vanish while giving rise to an NLE (robust boundary mode) localized around the hole. On the other hand, Fig. 4(b) and (d) demonstrate that introducing a hole defect by removing a single lattice site gives rise to new NLEs without altering the number of CLEs. Notably, in this case, the NLE (line mode) form connections between two holes rather than being confined to a single hole. To further explore the relationship between topological degeneracy and the number of hole defects, we calculate the number of zero-energy states, $N_{E=0}$, for varying numbers of holes, $N_{h}$, under the two distinct puncturing schemes. Using the Hamiltonian Eq. 1 with $\mu=0.1t$, we find that $N_{E=0}$ exhibits a linear dependence on $N_{h}$. Specifically, $N_{E=0}$ decreases as $N_{h}$ increases for the defect configuration shown in Fig. 4(a), while it grows with $N_{h}$ for the configuration depicted in Fig. 4(b). We also note that the number of zero-energy states remains unchanged when only a single hole defect present in the system for the configuration in Fig. 4(b). This is because the NLEs require two holes to form, as a single boundary cannot support linearly independent line excitations. In contrast, for boundary modes, the zero-energy degeneracy remains linearly dependent on the number of holes, even for a single hole. A similar phenomenon has recently been reported in the kagome lattice [75], where the flat-band degeneracy equals the sum of compact localized states and inner-robust boundary modes. Unlike the Lieb lattice, the kagome system lacks protected line modes due to the absence of boundary decoration. Our results establish a universal counting rule for singular flat-band degeneracy in holed systems:
@@ II.3 Lieb lattice with defects
III Entanglement property of Lieb Lattice
@@
Here we consider the Lieb lattice with periodic boundary condition as a quasi-one-dimensional system, as depicted in Fig. 5(a). Given that the van Hove singularities of the Lieb lattice are located at the K point, we enlarge the unit cell by $\sqrt{2}\times\sqrt{2}$. The Hamiltonian of the system is identical to Eq. 1, with the chemical potential set to zero for simplicity. First, we diagonalize the Hamiltonian to obtain its complete set of eigenstates and corresponding eigenenergies. As reflected in Fig. 5(b), two flat bands emerge from van Hove singularities at $E_{f}=\pm 2t$. Then we construct the correlation matrix $\Gamma_{A}$ for subsystem A, which is defined as follows:
@@
IV flat band from a perfect nesting van Hove system
Flat band structures can also arise from perfect nesting van Hove systems through a mechanism analogous to the one described earlier. The half-filled square lattice with nearest-neighbor hopping represents the simplest example of such a system, exhibiting perfect nesting and providing an ideal platform for engineering flat bands. The Fermi surface of the system comprises four straight lines connecting the four K points, which coincide with the van Hove points, as shown in Fig. 6(a). When periodic boundary conditions are applied in both the vertical and horizontal directions, the square lattice system supports loop excitation as illustrated in Fig. 6(b). However, if the vertical direction is treated as an open boundary while periodicity is preserved in the horizontal direction, two distinct scenarios emerge depending on the alignment of the top and bottom boundaries. In the matched boundary case, the number of layers between the top and bottom boundaries is odd, resulting in a configuration where the boundaries align consistently. On the other hand, the mismatched boundary case arises when the number of layers between the top and bottom boundaries is even, leading to a misalignment of the boundaries. These two configurations are depicted in Fig. 6(c) and (d), respectively. In the case of matched boundaries, line excitations connecting the top and bottom boundaries emerge as eigenstates of the Hamiltonian. In contrast, when the boundaries are mismatched, the wave function of the line excitations cannot undergo destructive interference at the top green points in Fig. 6(d). As a result, these excitations are no longer eigenstates of the system.
V conclusion and discussion
In this study, we demonstrate that embedding the Lieb lattice onto a topologically non-trivial manifold enables a direct mapping between non-contractible loop excitations and the generators of the manifold’s fundamental group. This connection demonstrates that singular flat band systems inherently exhibit topological degeneracy. We further show that the degeneracy of NLEs in the Lieb lattice depend on the specific hole configurations. Additionally, we explore the entanglement properties of a quasi-one-dimensional Lieb ring. The entanglement entropy displays jumps at the two van Hove singularities as the Fermi surface is varied. By systematically varying the system size, we observe that the entanglement entropy follows the characteristic behavior of a one-dimensional free massless complex fermion.
Since line modes in perfect nesting van Hove systems act as non-contractible loops of the manifolds, these systems offer another promising avenue for engineering flat bands. Using the square lattice as an example, we identify the specific boundary conditions required for the emergence of line modes in this system. Building on this insight, we design a square lattice system with strategically interspersed holes and compute its energy spectrum. The results reveal three flat bands overlapping at zero energy. Among these, two flat bands originate from two independent line excitations supported by each hole. The third band, however, is not intrinsic and is highly sensitive to the precise separation between adjacent holes. Recently, several articles have explored flat bands in other 2-dimensional systems with periodic holes [64, 65, 66, 67]. However, their approaches primarily involve replacing lattice sites with atomic substitutes, which differs significantly from our framework. In summary, these systems not only deepen our understanding of flat band formation in generic physical systems but also provide a versatile platform for engineering flat bands through precise geometric design.
VI ACKNOWLEDGMENTS
This work is supported by the Ministry of Science and Technology (Grant No. 2022YFA1403901), the National Natural Science Foundation of China (Grant No. NSFC-12494594, No. NSFC-11888101, No. NSFC-12174428), the Strategic Priority Research Program of the Chinese Academy of Sciences (Grant No. XDB28000000), the New Cornerstone Investigator Program, the Chinese Academy of Sciences through the Youth Innovation Promotion Association (Grant No. 2022YSBR-048) and the Shanghai Science and Technology Innovation Action Plan (Grant No. 24LZ1400800). We want to thank Prof Qian Niu, Yong-shi Wu, Yue Yu and Dr. Yuhao Gu for helpful discussion.

### v2

@@ Abstract
Introduction.— Topological order [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18] is a profound concept in condensed matter physics, describing phases of matter that cannot be characterized by traditional symmetry-breaking mechanisms [19]. Topological degeneracy is one of the most fundamental aspects of systems exhibiting topological order. Topological degeneracy refers to the ground state degeneracy of a system with topological order, which grows exponentially with the genus of the manifold on which the system is defined [20, 21, 22].
@@
In contrast to the toric code, systems with more complex topological order typically exhibit non-Abelian exchange phases [4, 8], indicating that their ground states are strongly-correlated and cannot be described within a quasi-particle framework. This strong correlation makes the construction of such topologically ordered systems on a lattice particularly challenging. However, in this work, we show that flat bands can emerge in systems exhibiting topological degeneracy without necessarily requiring non-Abelian exchange phases.
@@
In this Letter, we establish that embedding the Lieb lattice into a topologically non-trivial manifold creates a direct correspondence between non-contractible loop excitations and the generators of the fundamental group. This relationship reveals that singular flat band systems inherently exhibit topological degeneracy. Furthermore, we demonstrate that the degeneracy of flat bands in the Lieb lattice is determined by the specific hole configurations. Using the square lattice—a paradigmatic example of a perfectly nested van Hove system—we identify the boundary conditions necessary for the emergence of line modes. Building on this insight, we design a square lattice system with strategically placed holes and compute its energy spectrum. Our numerical results confirm the predicted appearance of zero-energy flat bands, which arise from line modes connecting pairs of holes.
Lieb lattice on a 1-genus surface.— The Hamiltonian of the Lieb lattice[68, 69, 76, 77], considering only nearest-neighbor hopping, is given by:
@@
Lieb lattice on a 2-genus surface.— As introduced previously, the linearly independent NLEs of a Lieb lattice with periodic boundary conditions can be geometrically interpreted as non-contractible loops of the torus. As a result, when the Lieb lattice is placed on a 2-genus surface, there are four linearly independent NLEs, as seen in Fig. 3(b). To validate this, we diagonalize the Hamiltonian in Eq. 1 with $\mu\neq 0$ to obtain all the zero-energy states, which collectively form the Hilbert space $V_{E=0}$. By determining the Hilbert space of all CLEs, denoted as $V_{CLE}$, the number of linearly independent NLEs, $N_{NLE}$, can be expressed as:
@@
Lieb lattice with defects.— The topological degeneracy of the Lieb lattice can also be manifested in an open-boundary Lieb lattice with multiple hole defects. Similar to the toric code [29], the Lieb lattice can be punctured in two physically distinct ways, leading to two different types of boundary conditions. These boundary conditions are crucial in shaping the spatial configurations of CLE and NLE. As illustrated in Fig. 4(a) and (c), removing five lattice sites to create a hole causes four CLEs to vanish while giving rise to an NLE (robust boundary mode) localized around the hole. On the other hand, Fig. 4(b) and (d) demonstrate that introducing a hole defect by removing a single lattice site gives rise to new NLEs without altering the number of CLEs. Notably, in this case, the NLE (line mode) form connections between two holes rather than being confined to a single hole. To further explore the relationship between topological degeneracy and the number of hole defects, we calculate the number of zero-energy states, $N_{E=0}$, for varying numbers of holes, $N_{h}$, under the two distinct puncturing schemes. Using the Hamiltonian Eq. 1 with $\mu=0.1t$, we find that $N_{E=0}$ exhibits a linear dependence on $N_{h}$. Specifically, $N_{E=0}$ decreases as $N_{h}$ increases for the defect configuration shown in Fig. 4(a), while it grows with $N_{h}$ for the configuration depicted in Fig. 4(b). We also note that the number of zero-energy states remains unchanged when only a single hole defect present in the system for the configuration in Fig. 4(b). This is because the NLEs require two holes to form, as a single boundary cannot support linearly independent line excitations. In contrast, for boundary modes, the zero-energy degeneracy remains linearly dependent on the number of holes, even for a single hole. A similar phenomenon has recently been reported in the kagome lattice [75], where the flat-band degeneracy equals the sum of compact localized states and inner-robust boundary modes. Unlike the Lieb lattice, the kagome system lacks protected line modes due to the absence of boundary decoration. Our results establish a universal counting rule for singular flat-band degeneracy in holed systems:
@@
Perfect nesting van Hove systems.— Flat band structures can also arise from perfect nesting van Hove systems through a mechanism analogous to the one described earlier. The half-filled square lattice with nearest-neighbor hopping represents the simplest example of such a system, exhibiting perfect nesting and providing an ideal platform for engineering flat bands. The Fermi surface of the system comprises four straight lines connecting the four K points, which coincide with the van Hove points, as shown in Fig. 5(a). When periodic boundary conditions are applied in both the vertical and horizontal directions, the square lattice system supports loop excitation as illustrated in Fig. 5(b). However, if the vertical direction is treated as an open boundary while periodicity is preserved in the horizontal direction, two distinct scenarios emerge depending on the alignment of the top and bottom boundaries. In the matched boundary case, the number of layers between the top and bottom boundaries is odd, resulting in a configuration where the boundaries align consistently. On the other hand, the mismatched boundary case arises when the number of layers between the top and bottom boundaries is even, leading to a misalignment of the boundaries. These two configurations are depicted in Fig. 5(c) and (d), respectively. In the case of matched boundaries, line excitations connecting the top and bottom boundaries emerge as eigenstates of the Hamiltonian. In contrast, when the boundaries are mismatched, the wave function of the line excitations cannot undergo destructive interference at the top green points in Fig. 5(d). As a result, these excitations are no longer eigenstates of the system.
Conclusion and discussion.— In this Letter, we demonstrate that placing the Lieb lattice on a topologically non-trivial manifold directly connects non-contractible loop excitations to the generators of the fundamental group, proving that singular flat band systems inherently host topological degeneracy. The flat band degeneracy depends critically on the lattice’s hole configurations. While this conclusion aligns with recent findings in Kagome systems [75], we have refined the specific formulas governing flat band degeneracy. Strikingly, quasi-1D Lieb rings exhibit entanglement entropy scaling governed by a free complex fermion conformal field theory, with sharp jumps at van Hove singularities as the Fermi surface varies (see Supplementary Material). We further derive the exact boundary conditions required for line modes in a square lattice—a paradigmatic example of a perfectly nested van Hove system. Inspired by this result, we engineer a square lattice with strategically placed holes, yielding a spectrum with three zero-energy flat bands. Two of these bands arise from independent line excitations and persist even when anisotropy breaks perfect nesting. Recently, several articles [64, 65, 66, 67] have explored flat bands in other 2-dimensional systems with periodic holes. However, their approaches primarily involve replacing lattice sites with atomic substitutes, which differs significantly from our framework. In summary, these systems not only deepen our understanding of flat band formation in generic physical systems but also provide a versatile platform for engineering flat bands through precise geometric design.
Acknowledgement.— This work is supported by the Ministry of Science and Technology (Grant No. 2022YFA1403901), the National Natural Science Foundation of China (Grant No. NSFC-12494594, No. NSFC-11888101, No. NSFC-12174428), the Strategic Priority Research Program of the Chinese Academy of Sciences (Grant No. XDB28000000), the New Cornerstone Investigator Program, the Chinese Academy of Sciences through the Youth Innovation Promotion Association (Grant No. 2022YSBR-048) and the Shanghai Science and Technology Innovation Action Plan (Grant No. 24LZ1400800). We want to thank Prof Qian Niu, Yong-shi Wu, Yue Yu and Dr. Yuhao Gu for helpful discussion.
Data availability. Numerical data for this manuscript are publicly accessible in Ref. [78].
Supplemental Material for ”Topological Degeneracy Induced Flat Bands in two-Dimensional Holed Systems”
Yuge Chen These three authors contributed equally
Hui Yu These three authors contributed equally
Yun-Peng Huang These three authors contributed equally
Zhen-Yu Zheng
Jiangping Hu
I Entanglement property of Lieb Lattice
@@
Here we consider the Lieb lattice with periodic boundary condition as a quasi-one-dimensional system, as depicted in Fig. 7(a). Given that the van Hove singularities of the Lieb lattice are located at the K point, we enlarge the unit cell by $\sqrt{2}\times\sqrt{2}$. The Hamiltonian of the system is identical to one descirbed in main text, with the chemical potential set to zero for simplicity. First, we diagonalize the Hamiltonian to obtain its complete set of eigenstates and corresponding eigenenergies. As reflected in Fig. 7(b), two flat bands emerge from van Hove singularities at $E_{f}=\pm 2t$. Then we construct the correlation matrix $\Gamma_{A}$ for subsystem A, which is defined as follows: